# Kähler compactification of $\mathbb{C}^{n}$ and Reeb dynamics

### v1

@@ 2.1. Fano cone singularities and AC Kähler metrics
    ([CH24]) Assume that the Reeb vector field ${\mathfrak{v}}$ is irregular. Then there exists a sequence of radius functions $\mathfrak{r}_{i}$ on $\mathcal{C}$ whose associated Reeb vector fields ${\mathfrak{v}}_{i}$ are quasi-regular such that ${\mathfrak{v}}_{i}$ converge to ${\mathfrak{v}}$ as $i\rightarrow+\infty$. For any $0<\epsilon\ll\lambda_{2}$ there exists $I$ such that for any $i\geq I$, $|\nabla_{g_{0,i}}^{j}(\Phi^{*}J-J_{0})|=O(\mathfrak{r}_{i}^{-\lambda_{2}+\epsilon-j})$ for all $j\in\mathbb{Z}_{\geq 0}$. As a consequence, there exists orbifold compactification of $W$ by $\mathcal{Y}_{i}=\mathcal{C}^{\circ}/\langle{\mathfrak{v}}_{i}\rangle$ when $i\geq I$.
[…]
@@ 2.2. Orbifold Kähler compactification of $\mathbb{C}^{n}$
We generalize some of Brenton-Morrow’s results to the orbifold setting. Let $(\mathbb{C}^{n},g)$ be an AC Kähler metric that is asymptotical to $(\mathcal{C},g_{0})$ with a quasi-regular Reeb vector field ${\mathfrak{v}}$. Let $\mathcal{Y}\subset\mathcal{X}$ be the orbifold compactification of $\mathbb{C}^{n}$ obtained by Theorem 2.4. In particular, $\mathcal{Y}=\mathcal{C}/\langle{\mathfrak{v}}\rangle$. Let $[\mathcal{Y}]$ denote the orbifold line bundle on $\mathcal{X}$ associated to the orbifold divisor $[\mathcal{Y}]$ and denote by $\mathcal{L}$ its restriction to $\mathcal{Y}$ which coincides with the orbifold normal bundle $\mathcal{L}=[\mathcal{Y}]|_{\mathcal{Y}}=N_{\mathcal{Y}/\mathcal{X}}$. We will use the orbifold cohomology introduced in [Hae84]. In other words, we set $H^{i}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})=H^{i}(B\mathcal{X},\mathbb{Z})$ and $H^{i}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=H^{i}_{\mathrm{orb}}(B\mathcal{Y},\mathbb{Z})$ where $B\mathcal{X}$ and $B\mathcal{Y}$ are the classifying spaces of the orbifolds $\mathcal{X}$ and $\mathcal{Y}$ respectively. $B\mathcal{X}$ can be constructed by the following procedure (see [Hae84, HS91] and [BG08, Chapter 4] for more details). Choose a Riemannian metric on $\mathcal{X}$ and let $P$ denote the differentiable bundle of orthonormal frames of $\mathcal{X}$. Let $EO(2n)\rightarrow BO(2n)$ be a universal $O(2n)$-bundle, then $B\mathcal{X}$ is the quotient of $EO(2n)\times P$ by the diagonal action of $O(2n)$. A similar construction applies to $\mathcal{Y}$. Moreover, since it is a global quotient of $M$ by an effective $S^{1}$ action we can use the following construction following [HS91, section 5.3]. We choose an $S^{1}$-invariant Riemannian metric on $M$. Let $P_{M}$ denote the bundle of orthonormal frames of $M$ and set $BM=P_{M}\times_{O(2n-1)}EO(2n-1)$ where $EO(2n-1)\rightarrow BO(2n-1)$ is a universal $O(2n-1)$-bundle. We can set $B\mathcal{Y}=BM/S^{1}$ and we have a principal $S^{1}$-bundle $BM\rightarrow BM/S^{1}=B\mathcal{Y}$.
@@
    $-K^{\mathrm{orb}}_{\mathcal{X}}=(r+1)[\mathcal{Y}]$ with $r\geq 1$ is an ample orbifold line bundle and $-K^{\mathrm{orb}}_{\mathcal{Y}}=r\mathcal{L}$ is also ample. In particular, both $\mathcal{X}$ and $\mathcal{Y}$ are Fano orbifolds and $(r+1)$, $r$ are their Fano indices respectively.
@@
Let $\overline{\mathcal{U}}$ be the closed neighborhood of $\mathcal{Y}\subset\mathcal{X}$ as obtained from Theorem 2.4. We denote by $\overline{U}$ the underlying space of orbifold $\overline{\mathcal{U}}$ (with smooth boundary). We have the exact sequence (for ordinary cohomology):
@@
Using excision property, $H^{k}(X-\overline{U},\partial\overline{U},\mathbb{Z})=H^{k}(X,\overline{U},\mathbb{Z})=H^{k}(X,Y,\mathbb{Z})=H^{k}_{c}(X\setminus Y,\mathbb{Z})$ for $0\leq k<2n$. So we have
@@
On the other hand, by Lefschetz duality, $H_{k}(X-\overline{U},\mathbb{Z})=H^{2n-k}(X-\overline{U},\partial\overline{U},\mathbb{Z})=0$ for $1\leq k\leq 2n$. By Universal Coefficients Theorem, $H^{k}(X-\overline{U},\mathbb{Z})=0$ for $1\leq k\leq 2n$. So we also get:
@@
We have $H^{k}_{\mathrm{orb}}(\mathcal{X},\mathcal{Y},\mathbb{Z})=H^{k}_{c}(\mathcal{X}\setminus\mathcal{Y},\mathbb{Z})=0$ for $0\leq k<2n$. So $H^{k}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=H^{k}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})$ for $20\leq k\leq 2n-$ 2. In particular $H^{0}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=H^{0}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})\cong\mathbb{Z}$ and hence $Y$ is connected. Since $H^{2n}_{\mathrm{orb}}(\mathcal{X},\mathcal{Y},\mathbb{Z})=H^{2n}_{c}(X\setminus Y,\mathbb{Z})=\mathbb{Z}$, when $k=2n-1$, we get:
(9) $0\rightarrow H^{2n-1}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})\rightarrow H^{2n-1}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})\rightarrow\mathbb{Z}{\rightarrow}H^{2n}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})\rightarrow H^{2n}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})\rightarrow 0$
For $k\geq 2n+1$, we get $H^{k}_{\mathrm{orb}}(\mathcal{X})\cong H^{k}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})$.
@@
where $\alpha_{k-1}$ is the cup product with $c_{1}(\mathcal{L})\in H^{2}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})$. By the vanishing property (7), we know that $\alpha_{k-1}$ is an isomorphism for $31\leq k\leq 2n-$ 3. Since $H^{0}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=\mathbb{Z}$, we get $H^{p}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=c_{1}^{\mathrm{orb}}(\mathcal{L})^{p/2}\mathbb{Z}$ for $20\leq p\leq 2n-$ 2 even. For $k=0$ in (10), we get $H^{1}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=0$ and we get $H^{p}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=0$ when $31\leq p\leq 2n-$ 3 is odd. By applying the same argument to $\mathbb{Q}$ coefficients, we also see that the orbifold cohomology ring of $\mathcal{Y}$ with $\mathbb{Q}$-coefficients is the same as the cohomology ring of $\mathbb{P}^{n-1}$ which is generated by $c_{1}(\mathcal{L})$.
@@
This implies in particular, $-K^{\mathrm{orb}}_{\mathcal{X}}=b[\mathcal{Y}]$ for some $b\in\mathbb{Z}$. If $b\leq 0$, then $K^{\mathrm{orb}}_{\mathcal{X}}$ has nonzero holomorphic section $s_{\mathcal{Y}}^{-b}$ which contradicts $\dim H^{0}_{\mathrm{orb}}(\mathcal{X},K^{\mathrm{orb}}_{\mathcal{X}})=h^{n,0}_{\mathrm{orb}}(X)=0$. So $b=1+r\geq 1$ i.e. $r\geq 0$. By adjunction formula, we get $-K^{\mathrm{orb}}_{\mathcal{Y}}=r\mathcal{L}$. If $r=0$, then $K^{\mathrm{orb}}_{\mathcal{Y}}$ is trivial and $H^{n-1}_{\mathrm{orb}}(\mathcal{Y},\mathcal{O}_{Y})=H^{0}_{\mathrm{orb}}(\mathcal{Y},K^{\mathrm{orb}}_{\mathcal{Y}})\cong\mathbb{C}$ which contradicts to the vanishing $h^{0,n-1}(\mathcal{Y})=0$. So $r\geq 1$.
By (7), we see that $M=\partial\bar{U}$ is an integral homology sphere. By the Mayer-Vietoris sequence to $\mathbb{C}^{n}=\overline{U}\cup(\mathbb{C}^{n}\setminus\overline{U})$, we know that $M=\overline{U}\cap(\mathbb{C}^{n}\setminus\overline{U})$ bounds an integral homology ball. ∎
@@
In general $H^{k}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})$ is not 0 when $k\geq 2n-1$. For example, it is well-known that for the weighted projective space $\mathbb{P}({\bf w})=\mathbb{P}(w_{1},w_{2},\dots,w_{n})$ with $d=\prod_{i}w_{i}$, we have the orbifold cohomology:
(12) $H^{k}_{\mathrm{orb}}(\mathbb{P}({\bf w}),\mathbb{Z})=\left\{\begin{array}[]{ll}\mathbb{Z}&k\leq 2n-2\text{ and is even}\\ \mathbb{Z}_{d}&k\geq 2n\text{ and is even}\\ 0&k\text{ is odd}.\end{array}\right.$
@@
    The Liouville vector field $X_{\lambda}$ defined by $\lambda=\iota_{X_{\lambda}}\mathrm{d}\lambda$ is transverse to $\partial W$ pointing outward/inward along $M_{+}/M_{-}$ respectively;
@@
$(W,\lambda)$ is called a Liouville filling of $(M_{+}.\xi_{+})$ if $M_{-}=\emptyset$.
If $M$ is the smooth boundary of strongly pseudo-convex domain $W$, then $(W,-\mathrm{d}^{\mathbb{C}}\rho)$ is a Liouville filling of $M$ whose contact structure is the CR structure, where $\rho$ is the strongly plurisubharmonic function whose sublevel let is $W$ with $M$ a regular level set. More generally, $(\rho^{-1}([a,b]),-\mathrm{d}^{\mathbb{C}}\rho)$ is a Liouville cobordism from $\rho^{-1}(a)$ to $\rho^{-1}(b)$ assuming both $a,b$ are regular values. In particular, by [H7̈3, Theorem 2.6.12], we have the following.
@@
The domain $W$ bounded by the contact link $M$ in $\mathbb{C}^{n}$ by 2.5 gives a Liouville filling of $M$.
In general, being a strongly pseudo-convex hypersurface in $\mathbb{C}^{n}$ does not imply the cobordism from $M$ to the standard $S^{2n-1}$ end of $\mathbb{C}^{n}$ is a Liouville cobordism. Thanks to the asymptotic cone structure here, we do have a Liouville cobordism in our case, which is crucial for our computation of symplectic homology.
@@
The cobordism from the contact link $M$ to the standard $S^{2n-1}$ end of $\mathbb{C}^{n}$ is a Liouville cobordism from the CR structure on $M$ to the standard contact/CR structure on $S^{2n-1}$.
@@
Since $M$, as a contact manifold, can be viewed as a level set of the radius function $\mathfrak{r}^{2}$, where $\mathfrak{r}^{2}$ defines a strongly plurisubharmonic function on the end of $\mathbb{C}^{n}$ by the asymptotic cone assumption. Let $\rho=|z|^{2}$ using the Euclidean metric, then $f(\rho)$ is plurisubharmonic on $\mathbb{C}^{n}$ if $f^{\prime\prime}\geq 0,f^{\prime}\geq 0$. We assume $M$ is contained in the ball $B(r)$ for $r\gg 0$. We can choose $f$, such that $f^{\prime\prime}\geq 0,f^{\prime}\geq 0$, $f(\rho)=0$ for $\rho\leq r^{2}$. Moreover, we require $f$ such that there exists $R\gg r$ so that $|\nabla\mathfrak{r}^{2}|\ll|\nabla f(\rho)|$ outside the ball $B(R)$. As a consequence, for $C\gg 0$, $C$ is a regular value of $F_{t}:=t\mathfrak{r}^{2}+f(\rho)$ for all $t\in[0,1]$. $F_{1}=\mathfrak{r}^{2}+f(\rho)$ is strongly plurisubharmonic outside $M$ and inside $F_{1}^{-1}(C)$, where $M$ is a regular level set of $F_{1}$ by construction. Therefore we have a Liouville cobordism from $M$ to $F_{1}^{-1}(C)$. Finally, as $F_{t}^{-1}(C)$ is a smooth family of strongly pseudo-convex hypersurfaces in $\mathbb{C}^{n}$, the Gray stability implies that $F_{1}^{-1}(C)=F_{0}^{-1}(C)$ as contact manifolds, while $F_{0}^{-1}(C)$ is the standard contact sphere by construction. ∎
@@
The vertex $o\in\mathcal{C}$ is a Gorenstein klt singularity.

### v3

@@ 2.1. Fano cone singularities and AC Kähler metrics
    ([CH24]) Assume that the Reeb vector field $\mathfrak{v}$} is irregular. Then there exists a sequence of radius functions $\mathfrak{r}_{i}$ on $\mathcal{C}$ whose associated Reeb vector fields ${\mathfrak{v}}_{i}$ are quasi-regular such that ${\mathfrak{v}}_{i}$ converge to $\mathfrak{v}$} inside the Reeb cone $\mathrm{t}_{\mathbb{R}}^{+}$ as $i\rightarrow+\infty$. For any $0<\epsilon\ll\lambda_{2}$ there exists $I$ such that for any $i\geq I$, $|\nabla_{g_{0,i}}^{j}(\Phi^{*}J-J_{0})|=O(\mathfrak{r}_{i}^{-\lambda_{2}+\epsilon-j})$ for all $j\in\mathbb{Z}_{\geq 0}$. As a consequence, there exists an orbifold compactification of $W$ by $\mathcal{Y}_{i}=\mathcal{C}^{\circ}/\langle{\mathfrak{v}}_{i}\rangle$ when $i\geq I$.
The compactification result in (1) was proved when $\mathfrak{v}$} is regular in [Li20] and generalized to the orbifold setting in [CH24, Appendix IV]. One way to get the diffeomorphism $\Phi$ is to use the deformation to the normal cone, which we now briefly explain. Assume first that $\mathfrak{v}$} is regular so that $\mathcal{Y}=(Y,\emptyset)$ and $\mathcal{X}=(X,\emptyset)$ in the above Theorem are smooth complex manifolds. Denote by $\mu:\mathfrak{X}\rightarrow X\times\mathbb{C}$ the blowing-up of $Y\times\{0\}\subset X\times\mathbb{C}$. The exceptional divisor $E$ is isomorphic to the projective bundle $\mathbb{P}(L\oplus\mathbb{C})$ where $L$ is the normal bundle of $Y$ in $X$. The divisor $E$ contains the divisor at the zero section $Y_{0}$ and the divisor $Y_{\infty}$ at infinity. We then get a flat family $\pi=\mathrm{proj}_{2}\circ\mu:\mathfrak{X}\rightarrow\mathbb{C}$ of projective varieties. Denote the fiber over $t\in\mathbb{C}$ by $X_{t}$. The central fiber $X_{0}$ is the union $X\cup E$ such that $X$ and $E$ are glued along $Y\subset X$ and $Y_{\infty}\subset E$. On a neighborhood of $Y_{0}\subset\mathfrak{X}$ we can integrate a smooth vector field to construct a smooth family of diffeomorphism $\sigma_{t}$ from a neighborhood $\mathcal{U}$ of $Y\subset L$ to a small neighborhood $\mathcal{V}_{t}$ of $Y\subset X_{t}$ such that $\sigma_{0}=\mathrm{id}_{\mathcal{U}}$ and $|\sigma_{t}^{*}J-J_{0}|=O(|t|)$ where $J$ is the complex structure on $X_{t}=X$ and $J_{0}$ is the complex structure on $L$ (see [Li20, Proposition 3.1]). Choose $\epsilon\ll 1$ such that $M^{\epsilon}=\{s\in L;|s|^{2}=\epsilon\}$ is contained in $\mathcal{U}$. It is clear that $M^{\epsilon}$ is isomorphic to $M$ as CR-manifolds thanks to the holomorphic $\mathbb{C}^{*}$ action on $L$. For $|t|\ll 1$, the CR manifolds $(M_{t},\sigma_{t}^{*}J|_{M^{\epsilon}})$ are still strongly pseudo-convex, whose associated contact structures are constant due to Gray’s stability theorem in contact geometry. More generally, if $\mathfrak{v}$} is quasi-regular, Conlon-Hein observed that the construction of deformation to the normal cone can be applied and moreover the restriction deformation of $\mathfrak{X}$ to a small neighborhood of $\mathcal{Y}\subset\mathcal{X}$ is locally trivial (see [CH24, Definition IV.5]). This can be considered as a tubular neighborhood theorem in the setting of cyclic orbifolds and allows us to use the same construction on uniformizing charts, like in the regular case, to get a diffeomorphism $\Phi$ as stated in the above theorem.
@@ 2.2. Orbifold Kähler compactification of $\mathbb{C}^{n}$
First, we recall some results proved by Brenton-Morrow in [BM78] (see also [vdV62, PS91]).
Theorem 2.8.
[BM78, Theorem 1.1] Let $X$ be a connected compact complex manifold and let $Y$ be a Kähler submanifold such that $X\setminus Y$ is biholomorphic to $\mathbb{C}^{n}$. Then the following statements are true:
  1. (1)
    The complex manifold $X$ is projective algebraic and $Y$ is positively embedded hypersurface. The line bundle $[Y]$ associated to the divisor $Y$ is ample over $X$.
  2. (2)
    There are ring isomorphisms $H^{*}(X,\mathbb{Z})\cong H^{*}(\mathbb{P}^{n},\mathbb{Z})$ and $H^{*}(Y,\mathbb{Z})\cong H^{*}(\mathbb{P}^{n-1},\mathbb{Z})$.
  3. (3)
    $H^{2}(X,\mathbb{Z})$ is generated by $c_{1}([Y])$ and $H^{2}(Y,\mathbb{Z})$ is generated by $c_{1}(L)$ where $L=[Y]|_{Y}$ is the normal bundle of $Y$ inside $X$.
  4. (4)
    $K_{X}=(r+1)[Y]$ with $r\geq 1$ and $K_{Y}=rL$. In particular, $X$ and $Y$ are both Fano manifolds.
Remark 2.9.
Rigorously speaking, in [BM78], the above results were proved in a more general situation that $X\setminus Y$ is a complex homology cell, and for that purpose, another assumption was added: $X$ contains no exceptional subvarieties. This assumption was used in the proof first to show that the normal bundle $L=[Y]|_{Y}\rightarrow Y$ is positive (it was shown that $L$ is either positive or negative) and then to show that $X$ is projective via a result of Grauert ([Gra62], [MR75, Theorem 2.4]). Here we already know that $X\setminus Y$ is $\mathbb{C}^{n}$, which is Stein, hence it is guaranteed that the normal bundle $L$ is positive (since otherwise by Grauert’s criterion ([Gra62]) $Y$ can be contracted to a point $p$ to give a compact analytic space $\hat{X}$, but $X\setminus Y=\hat{X}\setminus p$ can never be Stein when the dimension $n>1$).
Now we generalize the above results to an orbifold setting. Let $(\mathbb{C}^{n},g)$ be an AC Kähler metric that is asymptotical to $(\mathcal{C},g_{0})$ with a quasi-regular Reeb vector field $\mathfrak{v}$}. Let $\mathcal{Y}\subset\mathcal{X}$ be the orbifold compactification of $\mathbb{C}^{n}$ obtained by Theorem 2.7. In particular, $\mathcal{Y}=\mathcal{C}^{\circ}/\langle{\mathfrak{v}}\rangle$. Let $\mathcal{Y}]$ denote the orbifold line bundle on $\mathcal{X}$ associated to the orbifold divisor $\mathcal{Y}]$ and denote by $\mathcal{L}$ its restriction to $\mathcal{Y}$ which coincides with the orbifold normal bundle $\mathcal{L}=[\mathcal{Y}]|_{\mathcal{Y}}=N_{\mathcal{Y}/\mathcal{X}}$. We refer to [BG08, 4.4] for these notions for complex orbifolds. We will use the orbifold cohomology introduced in [Hae84]. In other words, we set $H^{i}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})=H^{i}(B\mathcal{X},\mathbb{Z})$ and $H^{i}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=H^{i}_{\mathrm{orb}}(B\mathcal{Y},\mathbb{Z})$ where $B\mathcal{X}$ and $B\mathcal{Y}$ are the classifying spaces of the orbifolds $\mathcal{X}$ and $\mathcal{Y}$ respectively. $B\mathcal{X}$ can be constructed by the following procedure (see [Hae84, HS91] and [BG08, Chapter 4] for more details). Choose a Riemannian metric on $\mathcal{X}$ and let $P$ denote the differentiable bundle of orthonormal frames of $\mathcal{X}$. Let $EO(2n)\rightarrow BO(2n)$ be a universal $O(2n)$-bundle, then $B\mathcal{X}$ is the quotient of $EO(2n)\times P$ by the diagonal action of $O(2n)$. A similar construction applies to $\mathcal{Y}$. Moreover, since it is a global quotient of $M$ by an effective $S^{1}$ action, we can use the following construction following [HS91, section 5.3]. We choose an $S^{1}$-invariant Riemannian metric on $M$. Let $P_{M}$ denote the bundle of orthonormal frames of $M$ and set $BM=P_{M}\times_{O(2n-1)}EO(2n-1)$ where $EO(2n-1)\rightarrow BO(2n-1)$ is a universal O(2n1)O(2n-1)-bundle. We can set $B\mathcal{Y}=BM/S^{1}$ and we have a principal $S^{1}$-bundle $BM\rightarrow BM/S^{1}=B\mathcal{Y}$.
@@
    $K^{\mathrm{orb}}_{\mathcal{X}}=(r+1)[\mathcal{Y}]$ with $r\geq 1$ is an ample orbifold line bundle and $K^{\mathrm{orb}}_{\mathcal{Y}}=r\mathcal{L}$ is also ample. In particular, both $\mathcal{X}$ and $\mathcal{Y}$ are Fano orbifolds and $(r+1)$, $r$ are their Fano indices respectively.
@@
Let $\overline{\mathcal{U}}$ be the closed neighborhood of $\mathcal{Y}\subset\mathcal{X}$ as obtained from Theorem 2.7. We denote by $\overline{U}$ the underlying space of the orbifold $\overline{\mathcal{U}}$ (with smooth boundary). We have an exact sequence for ordinary cohomology:
@@
Using the excision property, we have $H^{k}(X-\overline{U},\partial\overline{U},\mathbb{Z})=H^{k}(X,\overline{U},\mathbb{Z})=H^{k}(X,Y,\mathbb{Z})=H^{k}_{c}(X\setminus Y,\mathbb{Z})$ for $0\leq k<2n$. So we have
@@
On the other hand, by the Lefschetz duality, we have $H_{k}(X-\overline{U},\mathbb{Z})=H^{2n-k}(X-\overline{U},\partial\overline{U},\mathbb{Z})=0$ for $1\leq k\leq 2n$. By the Universal Coefficients Theorem, $H^{k}(X-\overline{U},\mathbb{Z})=0$ for $1\leq k\leq 2n$. So we also get:
@@
We have $H^{k}_{\mathrm{orb}}(\mathcal{X},\mathcal{Y},\mathbb{Z})=H^{k}_{c}(\mathcal{X}\setminus\mathcal{Y},\mathbb{Z})=0$ for $0\leq k<2n$. Therefore we get $H^{k}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=H^{k}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})$ for $0\leq k\leq 2n-2$. In particular $H^{0}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=H^{0}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})\cong\mathbb{Z}$ and hence $Y$ is connected. Since $H^{2n}_{\mathrm{orb}}(\mathcal{X},\mathcal{Y},\mathbb{Z})=H^{2n}_{c}(X\setminus Y,\mathbb{Z})=\mathbb{Z}$ when k=2n1k=2n-1, we get:
(10) $0\rightarrow H^{2n-1}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})\rightarrow H^{2n-1}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})\rightarrow\mathbb{Z}{\rightarrow}H^{2n}_{\mathrm{orb}}(\mathcal{X},\mathbb{Z})\rightarrow H^{2n}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})\rightarrow 0.$
Hence for $k\geq 2n+1$, we get $H^{k}_{\mathrm{orb}}(\mathcal{X})\cong H^{k}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})$.
@@
where $\alpha_{k-1}$ is the cup product with $c_{1}(\mathcal{L})\in H^{2}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})$. By the vanishing property (8), we know that $\alpha_{k-1}$ is an isomorphism for $1\leq k\leq 2n-3$. Since $H^{0}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=\mathbb{Z}$, we get $H^{p}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=c_{1}^{\mathrm{orb}}(\mathcal{L})^{p/2}\mathbb{Z}$ for $0\leq p\leq 2n-2$ even. For $k=0$ in (11), we get $H^{1}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=0$ and we get $H^{p}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})=0$ when $1\leq p\leq 2n-3$ is odd. By applying the same argument to $\mathbb{Q}$ coefficients, we also see that the orbifold cohomology ring of $\mathcal{Y}$ with $\mathbb{Q}$-coefficients is the same as the cohomology ring of $\mathbb{P}^{n-1}$, which is generated by $c_{1}(\mathcal{L})$.
@@
This implies in particular, $K^{\mathrm{orb}}_{\mathcal{X}}=b[\mathcal{Y}]$ for some $b\in\mathbb{Z}$. If $b\leq 0$, then $K^{\mathrm{orb}}_{\mathcal{X}}$ has nonzero holomorphic section $s_{\mathcal{Y}}^{-b}$ which contradicts $\dim H^{0}_{\mathrm{orb}}(\mathcal{X},K^{\mathrm{orb}}_{\mathcal{X}})=h^{n,0}_{\mathrm{orb}}(X)=0$. So $b=1+r\geq 1$ i.e. $r\geq 0$. By the adjunction formula, we get $K^{\mathrm{orb}}_{\mathcal{Y}}=r\mathcal{L}$. If $r=0$, then $K^{\mathrm{orb}}_{\mathcal{Y}}$ is trivial and $H^{n-1}_{\mathrm{orb}}(\mathcal{Y},\mathcal{O}_{Y})=H^{0}_{\mathrm{orb}}(\mathcal{Y},K^{\mathrm{orb}}_{\mathcal{Y}})\cong\mathbb{C}$ which contradicts to the vanishing $h^{0,n-1}(\mathcal{Y})=0$. So $r\geq 1$.
By (8), we see that $M=\partial\bar{U}$ is an integral homology sphere. By applying the Mayer-Vietoris sequence to $\mathbb{C}^{n}=(\overline{U}\backslash\mathcal{Y})\cup(\mathbb{C}^{n}\setminus U)$, we know that $M=\overline{U}\cap(\mathbb{C}^{n}\setminus U)$ bounds an integral homology ball $\mathbb{C}^{n}\setminus U$. ∎
@@
In general, $H^{k}_{\mathrm{orb}}(\mathcal{Y},\mathbb{Z})$ is not 0 when $k\geq 2n-1$. For example, it is well-known that for the weighted projective space $\mathbb{P}({\bf w})=\mathbb{P}(w_{1},w_{2},\dots,w_{n})$ with $d=\prod_{i}w_{i}$, we have the orbifold cohomology:
[…]
We will now show that the (2n1)(2n-1)-dimensional submanifold $M$ of $\mathbb{C}^{n}$ from the above proposition, equipped with the induced contact structure, is cobordant to the standard contact sphere via a Liouville cobordism.
@@
    The Liouville vector field $X_{\lambda}$ defined by $\lambda=\iota_{X_{\lambda}}\mathrm{d}\lambda$ is transversal to $\partial W$ pointing outward/inward along $M_{+}/M_{-}$ respectively;
@@
$(W,\lambda)$ is called a Liouville filling of $(M_{+},\xi_{+})$ if $M_{-}=\emptyset$.
If $M$ is the smooth boundary of strongly pseudo-convex domain $W$, then $(W,\mathrm{d}^{C}\rho:=-Jd\rho)$ is a Liouville filling [^3] Note that in [Sei08, (3.15)], $(\mathrm{d}^{C}\rho)(v)=\mathrm{d}\rho(Jv)=(J\mathrm{d}\rho)(v)$. This explains the potential sign difference in some symplectic literature. of $M$ whose contact structure is induced by the CR structure and $\rho$ is the strongly plurisubharmonic function whose sub-level set is $W$ with $M$ a regular level set. More generally, $(\rho^{-1}([a,b]),\mathrm{d}^{C}\rho)$ is a Liouville cobordism from $\rho^{-1}(a)$ to $\rho^{-1}(b)$ assuming both $a,b$ are regular values. In particular, by [H7̈3, Theorem 2.6.12], we have the following.
@@
The domain $W$ bounded by the contact link $M$ in $\mathbb{C}^{n}$ by 2.10 gives a Liouville filling of $M$.
In general, being a strongly pseudo-convex hypersurface in $\mathbb{C}^{n}$ does not imply that the cobordism from $M$ to the standard $S^{2n-1}$ end of $\mathbb{C}^{n}$ is a Liouville cobordism. Thanks to the asymptotic cone structure, we do have a Liouville cobordism in our case, which is crucial for our computation of symplectic homology.
@@
The cobordism from the contact link $M$ to the standard $S^{2n-1}$ end of $\mathbb{C}^{n}$ admits a Liouville cobordism structure from the CR structure on $M$ to the standard contact/CR structure on $S^{2n-1}$.
@@
Since $M$, as a contact manifold, can be viewed as a level set of the radius function $\mathfrak{r}^{2}$, where $\mathfrak{r}^{2}$ defines a strongly plurisubharmonic function on the end of $\mathbb{C}^{n}$ by the asymptotic cone assumption. Let $\rho=|z|^{2}$ using the Euclidean metric. Then $f(\rho)$ is plurisubharmonic on $\mathbb{C}^{n}$ if $f^{\prime\prime}\geq 0,f^{\prime}\geq 0$. We assume $M$ is contained in the ball $B(r)$ for a fixed $r\gg 1$. We can choose $f$, such that $f^{\prime\prime}\geq 0,f^{\prime}\geq 0$, $f(\rho)=0$ for $\rho\leq r^{2}$. Fix $R>r$ and we require $f$ to grow sufficient fast such that so that $|\nabla\mathfrak{r}^{2}|<\frac{1}{2}|\nabla f(\rho)|$ outside the ball $B(R)$. As a consequence, for some $\gg 1, CC$ is a regular value of $F_{t}:=t\mathfrak{r}^{2}+f(\rho)$ for all $t\in[0,1]$. The function $F_{1}=\mathfrak{r}^{2}+f(\rho)$ is strongly plurisubharmonic outside $M$ and inside $F_{1}^{-1}(C)$, where $M$ is a regular level set of $F_{1}$ by construction. Therefore, we have a Liouville cobordism from $M$ to $F_{1}^{-1}(C)$. Finally, as $F_{t}^{-1}(C)$ is a smooth family of strongly pseudo-convex hypersurfaces in $\mathbb{C}^{n}$, the Gray stability implies that $F_{1}^{-1}(C)$ is isomorphic to $F_{0}^{-1}(C)$ as contact manifolds, while $F_{0}^{-1}(C)$ is the standard contact sphere by construction. ∎
For the rest of this subsection, we study the algebraic property of the singularity $(\mathcal{C},o)$, which is again the asymptotical cone associated to a complete AC Kähler metric on $(\mathbb{C}^{n},J_{0})$ obtained by Theorem 2.7.
@@
With the above notation, the vertex $o\in\mathcal{C}$ is a Gorenstein klt singularity.